# Sublogarithmic Distributed Vertex Coloring with Optimal Number of Colors

### 30 Mar 2026 v1

@@ 1. $E_{CC}(i)$:
Step (S1) holds with a similar reasoning as in 5.3 (and because $F$ includes the cliques) for $c_{1}=5$. Step (S2) holds for $c_{2}=2$ as a variable is only contained in events with variables in its $2$-hop neighborhood. Step (S3) follows directly with the definition of $\mathcal{A}^{\prime}$. Step (S4) requires an upper bound of $1/\Delta(F)^{c_{4}}$ on the probability for each bad event $A\in\mathcal{A}$ to be contained in $\mathcal{A}^{\prime}$ where $c_{1}$ is an arbitrary constant satisfying $c_{4}>3c_{1}(4c_{\mathcal{B}^{\prime}}+16)+c_{2}+c_{\mathcal{B}^{\prime}}+1$. By claim 6.11 this probability is at most $\Delta(G)\cdot\exp(-\Delta(G)^{1/40})$. As $\Delta(G)=\Theta(\Delta(F))$ Step (S4) holds for $c_{4}$ (as chosen above) as long as $\Delta(G)$ is larger than a sufficiently large absolute constant $\Delta_{0}$. The set of events included in the LLL of Algorithm 8 are all included in $\mathcal{B}^{\prime}$ with the $c_{\mathcal{B}^{\prime}}$ defined above; a smaller choice of $c_{\mathcal{B}^{\prime}}$ is possible, but it is more tedious to verify that all events are then included in $\mathcal{B}^{\prime}$. Hence by Lemma 3.2, the connected components of $G_{\operatorname{dep}}$ have size at most $O(\Delta(F)^{4c_{1}c_{\mathcal{B}^{\prime}}+4c_{1}+c_{2}}\log n)=\operatorname{\text{{\rm poly}}}(\log n)$ with high probability.
@@
It remains to verify that the claimed properties indeed hold. CC is respected during both pre-shattering and post-shattering so Part (b) holds as well.
@@ 7 Coloring Cliques
The goal of this section is to extend the coloring to the cliques in $A_{H}$ and $A_{L}$, respectively. In Algorithm 1, the cliques of $A_{H}$ are colored first, and thus when coloring these nodes we need to ensure that CC is satisfied for all cliques in $A_{L}$. We prove the following:
@@
At a high level, we follow the outline of the approach of [45]. In the first phase, we compute a defective coloring (i.e., a coloring with monochromatic edges) of the cliques to be colored by assigning a random permutation of the colors not used in $\mathrm{All}_{i}$ to the vertices of $A_{i}$. About $\sqrt{\Delta}$ vertices of each clique have a conflict in that they are assigned the same color as an external neighbor (see Lemma 7.2(a)). In the second phase, we perform color swaps: each node with a conflict finds a partner within its own clique with whom it can swap colors to remove the conflict without introducing new ones. This is where the CC property is crucially used.
@@ 7.1 Step 1: Synchronized Color Trial
    All cliques of $A^{\prime}$ and all uncolored cliques not in $A^{\prime}$ satisfy the CC property.
@@ 7.2 Step 2: Finding Safe Swaps
We next show that a constant fraction of the nodes of $A_{i}$ are swappable (for any given node $v$). This lemma is the only place where we use the CC property. Its proof is purely deterministic, given the CC property of the clique and the guarantees of Lemma 7.2.
@@
Nodes violating (c): Note that CC holds for the clique $A_{i}$, as it held before Step 1 (recall Lemma 4.8(1)) and was maintained by Step 2 (by Lemma 7.2(b)). Hence, by 4.3 each color (here applied to $\gamma(v)$) appears on an external neighbor (recall, one outside of $A_{i}\cup\mathrm{All}_{i}$) of at most $4\Delta/5$ vertices in $A_{i}$.
@@ 3. $iii)$
    (Strong version of CC) For each uncolored $A_{j}\not\in A^{\prime}$ and color $x$, at most $2\cdot\Delta^{37/40}$ vertices in $A_{j}$ have a neighbor outside of $A_{j}\cup\mathrm{All}_{j}\cup\operatorname{Big}_{j}^{+}$ that has a candidate of color $x$ or is a candidate for a node of color $x$.
@@ 7.3 Step 3: Local Matching
The remaining task – deciding on swap partners – then reduces to a local bipartite matching problem. Consider the bipartite graph with vertex bipartition $L=\mathrm{Unhappy}_{i}$ and $R=A_{i}\setminus\mathrm{Unhappy}_{i}$, with an edge between each $v\in L$ and $u\in R$ if and only if $u\in T_{v}$. By Lemma 7.6(1,2), each node in $L$ has at least $\Delta^{17/40}/40$ incident edges, while each node in $R$ has at most $\Delta^{17/40}/80$ incident edges. It follows that for every subset $S\subseteq L$, $|N(S)|\geqslant|S|$. Then, by Hall’s theorem there is a matching $M$ that saturates $L$. We perform exactly the color swaps given by the matching $M$, which yields a proper coloring by 7.7.
To conclude with the proof of Lemma 4.8, it remains to argue that the CC property is maintained for the remaining uncolored cliques. This follows for Step 1 by Lemma 7.2(b) and for Step 3 by Lemma 7.6(3).
@@ 2. $E_{b}(i)$
    : $A_{i}$ is adjacent to a clique of $A^{\prime}$ or to an uncolored clique not in $A^{\prime}$ for which the CC property was broken.
@@
Let us now bound ${\mathbb{P}}[E_{b}(i)]$. Fix an arbitrary color $x$ and a clique $A_{j}$. Due to Lemma 4.1 the external degree of $A_{i}$ is bounded by $10^{8}\sqrt{\Delta}$. For a fixed external neighbor $u\in A_{\ell}\in A^{\prime\prime}$ of $A_{j}$, the probability that it receives color $x$ in the random permutation of $A_{j}$ is $1/|A_{j}|\leqslant 2/\Delta$ (by Lemma 4.1(b)). Furthermore, at most one vertex in each clique is assigned color $x$ and the random permutation for different cliques are independent. Hence, we can apply Lemma 4.4 with $Q=10^{8}\sqrt{\Delta}$ as it bounds external degrees from above (by Lemma 4.1(3,5)), and get that $A_{j}$ does not satisfy CC for color $x$ with probability at most $\exp(-\Delta^{1/40})$. By union bound on all colors and cliques adjacent to $A_{i}$, the event $E_{b}(i)$ occurs w.p. at most $\Delta^{3/2+1}\cdot\exp(-\Delta^{1/40})\leqslant\exp(-\Omega(\Delta^{1/40}))$.
@@
Consider the variable graph $H$ of this random process: it has a vertex (i.e., a variable) $v_{i}$ for each $A_{i}\in A^{\prime}$ and vertices $v_{i}$ and $v_{j}$ are connected iff an edge connects $A_{i}$ and $A_{j}$. The collections $\mathcal{A}$ and $\mathcal{B}$ from Lemma 3.2 are all 1-hop neighborhoods in $H$, i.e., all the $\{v_{i}\}\cup N_{H}(v_{i})$ for $A_{i}\in A^{\prime}$. Define $c_{\mathcal{B}^{\prime}}=0$ so that the set $\mathcal{B}^{\prime}$ in Lemma 3.2 includes the 1-hop neighborhood of every $v_{i}$ with $A_{i}\in A^{\prime\prime}$. This is a superset of $A^{\prime\prime}$ (the set of cliques for which we have an event in post-shattering), so it suffices to bound from above the sizes of the connected components induced by the events corresponding to sets of $\mathcal{B}^{\prime}$. Steps (S1) and (S2) with $c_{1}=2$ and $c_{2}=1$ are direct from the choice of sets $\{v_{i}\}\cup N_{H}(v_{i})$ because it has diameter at most 2 and every $v_{i}$ is included only in the sets of its neighbors. Step (S3) is direct from Algorithm 9. By Claim 7.8, each $A_{i}$ is included in $A^{\prime\prime}$ (i.e., a bad event occurred and $A_{i}$ uncolored its vertices) with probability at most $\exp(-\Omega(\Delta^{1/40}))$. The degree of a $v_{i}$ is at most $10^{9}\Delta^{3/2}$, because every vertex in $A_{i}$ may be adjacent to up to $10^{9}\sqrt{\Delta}$ vertices in other cliques. It implies that for $\Delta$ sufficiently large, we can pick a constant $c_{4}>3c_{1}(4c_{\mathcal{B}^{\prime}}+16)+c_{2}+c_{\mathcal{B}^{\prime}}+1=O(1)$ such that each set of $\mathcal{A}$ is kept in $\mathcal{A}^{\prime}$ with probability at most $\Delta(H)^{-c_{4}}$, i.e., Step (S4). Therefore Lemma 3.2 applies and, w.h.p., we can bound the size of the largest component in the dependency graph of the LLL of Algorithm 10 by some $\operatorname{\text{{\rm poly}}}(\log n)$. Let us conclude the proof by proving that the resulting coloring indeed satisfies (a) and (b). For (b), it simply follows from the fact that after retractions (Algorithm 10), the CC is preserved, and since events $E_{b}(i)$ are avoided in post-shattering, CC is also preserved by this coloring step (with a fresh budget). To see why (a) holds, observe that every $A\notin A^{\prime\prime}$ satisfies (a) after Algorithm 10 (otherwise $E_{a}(i)$ occurs and $A_{i}$ is included in $A^{\prime\prime}$). Importantly, if a monochromatic edge $\{u,v\}$ appears where $v$ was colored by Algorithm 10 and $u$ belongs to a clique of $A^{\prime}$ that did not participate in post-shattering (not in $A^{\prime\prime}$), we orient the edge from $u$ to $v$. So, the coloring of a clique $A_{j}$ that participates in the post-shattering phase can never increase the size of $\mathrm{Unhappy}_{i}$ of an adjacent clique $A_{i}$ that does not participate in the post-shattering phase. Finally, a clique $A_{i}\in A^{\prime\prime}$ colored by Algorithm 10 satisfies (a) at the end of the algorithm, as the event $E_{a}(i)$ would otherwise occur in the post-shattering LLL.
@@ 5. $B_{4}(i)$:
Fix a clique $A_{j}$ (either in $A^{\prime}$ or uncolored) and a color $x$. For a fixed external neighbor $u\in A_{\ell}\neq A_{j}$, the vertex with color $x$ in $u$’s clique (if any) gets sampled into $u$’s candidate set $S_{u}$ or samples $u$ into its candidate set with probability at most $p\leqslant\Delta^{-23/40}$. On the other hand, by assumption in Lemma 4.8, the external degrees of uncolored cliques not in $A^{\prime}$ is at most $30\Delta^{1/4}$ (recall that in this step, we do not need to guarantee that CC is maintained by cliques of $A^{\prime}$, in particular, cliques of $A_{H}$ do not need to verify Lemma 7.6(3)). And since the sampling is independent, it is easy to verify that (P7.1) holds with $Q=30\Delta^{1/4}$, thus that $B_{3}(i)$ occurs because of clique $A_{j}$ with probability at most $\exp(-\Delta^{1/40})$ by Lemma 4.4. By union bound on the colors and the at most $10^{8}\Delta^{3/2}$ cliques adjacent to $A_{i}$, we obtain the desired probability bound.
@@ 8 Ultrafast Coloring High-Degree Graphs
If it was not for the CC constraint, one could simply run the $(\deg+1)$-list-coloring algorithm of [35] to extend the coloring to $H$. Given an instance of $(\deg+1)$-list-coloring, it is not possible in general to ensure that the CC constraint will be maintained. Here, we use that, thanks to $\Pi$, the nodes of $H$ have lists significantly larger than their degree.
@@
Using $\Pi$ and Lemma 4.4, we can show that each of the $O(\log^{*}n)$ steps of [35] maintain the CC property with high probability.
@@
Both algorithms begin with a random color trial — for the sparse vertices, only some vertices try a random color. See [35, Algorithm 3]. Since every vertex of $H$ samples one color out of a list of $U\Delta^{0.22}$ while the external degree of uncolored cliques is upper bounded by $U$ (by $\Pi$), Lemma 4.4 and union bound implies that CC is not maintained for some clique with probability at most $n\Delta\exp(-\Delta^{1/40})<1/\operatorname{\text{{\rm poly}}}(n)$. Note that the dummy nodes do not verify $\Pi$, as their lists are of size two, but they are not adjacent to the cliques of $F$.
@@
Hence, we can apply Lemma 4.4 as before to reason that each step of MCT maintains CC with probability at least $1-\exp(-\Omega(\sqrt{s_{\min}}))-\Delta\exp(-\Omega(s_{\min}))\geqslant 1-1/\operatorname{\text{{\rm poly}}}(n)$.
The analysis of slack color for nodes of $H[C_{1}\cup\ldots\cup C_{q}]$ is the same as above. We emphasize that we do not need the put-aside set, Step 3 and 7 in [35, Algorithm 5], because $\Pi$ ensures that all nodes of $H$ have more than $\log^{2}n$ colors in their lists. After slack generation, [35, Algorithm 5] runs a synchronized color trial before it calls SlackColor on (most) nodes. See [35, Algorithm 7]. In this algorithm, (most of) the vertices in each $C_{i}$ receive a random color from a selected leader. As such, the probability that any given color is tried by a vertex in some $C_{i}$ is $O(1/|C_{i}|)$. Since the vertices have $\deg(v)=|L(v)|-1\geqslant U\Delta^{0.22}$, where the equality holds because of the dummy vertices and the inequality because of $\Pi$, an almost-clique $C_{i}$ contains at least $\Omega(U\Delta^{0.22})$ vertices. And thus the CC constraint is maintained by the SCT as every vertex tries gets some color $x$ with probability at most $O(1/|C_{i}|)\leqslant O(\frac{1}{U\Delta^{0.22}})\leqslant\frac{1}{U\Delta^{1/5}}$.

### 01 Apr 2026 v2

@@ 1. $E_{CC}(i)$:
Step (S1) holds with a similar reasoning as in 5.3 (and because $F$ includes the cliques) for $c_{1}=5$. Step (S2) holds for $c_{2}=2$ as a variable is only contained in events with variables in its 22-hop neighborhood. Step (S3) follows directly with the definition of $\mathcal{A}^{\prime}$. Step (S4) requires an upper bound of $1/\Delta(F)^{c_{4}}$ on the probability for each bad event $A\in\mathcal{A}$ to be contained in $\mathcal{A}^{\prime}$ where $c_{1}$ is an arbitrary constant satisfying $c_{4}>3c_{1}(4c_{\mathcal{B}^{\prime}}+16)+c_{2}+c_{\mathcal{B}^{\prime}}+1$. By claim 6.11 this probability is at most $\Delta(G)\cdot\exp(-\Delta(G)^{1/40})$. As $\Delta(G)=\Theta(\Delta(F))$ Step (S4) holds for $c_{4}$ (as chosen above) as long as $\Delta(G)$ is larger than a sufficiently large absolute constant $\Delta_{0}$. The set of events included in the LLL of Algorithm 8 are all included in $\mathcal{B}^{\prime}$ with the $c_{\mathcal{B}^{\prime}}$ defined above; a smaller choice of $c_{\mathcal{B}^{\prime}}$ is possible, but it is more tedious to verify that all events are then included in $\mathcal{B}^{\prime}$. Hence by Lemma 3.2, the connected components of $G_{\operatorname{dep}}$ have size at most $O(\Delta(F)^{4c_{1}c_{\mathcal{B}^{\prime}}+4c_{1}+c_{2}}\log n)=\operatorname{\text{{\rm poly}}}(\log n)$ with high probability.
@@
It remains to verify that the claimed properties indeed hold. $C$ is respected during both pre-shattering and post-shattering so Part (b) holds as well.
@@ 7 Coloring Cliques
The goal of this section is to extend the coloring to the cliques in $A_{H}$ and $A_{L}$, respectively. In Algorithm 1, the cliques of $A_{H}$ are colored first, and thus when coloring these nodes we need to ensure that $C$ is satisfied for all cliques in $A_{L}$. We prove the following:
@@
At a high level, we follow the outline of the approach of [45]. In the first phase, we compute a defective coloring (i.e., a coloring with monochromatic edges) of the cliques to be colored by assigning a random permutation of the colors not used in $\mathrm{All}_{i}$ to the vertices of $A_{i}$. About $\sqrt{\Delta}$ vertices of each clique have a conflict in that they are assigned the same color as an external neighbor (see Lemma 7.2(a)). In the second phase, we perform color swaps: each node with a conflict finds a partner within its own clique with whom it can swap colors to remove the conflict without introducing new ones. This is where the $C$ property is crucially used.
@@ 7.1 Step 1: Synchronized Color Trial
    All cliques of $A^{\prime}$ and all uncolored cliques not in $A^{\prime}$ satisfy the $C$ property.
@@ 7.2 Step 2: Finding Safe Swaps
We next show that a constant fraction of the nodes of $A_{i}$ are swappable (for any given node $v$). This lemma is the only place where we use the $C$ property. Its proof is purely deterministic, given the $C$ property of the clique and the guarantees of Lemma 7.2.
@@
Nodes violating (c): Note that $C$ holds for the clique $A_{i}$, as it held before Step 1 (recall Lemma 4.8(1)) and was maintained by Step 2 (by Lemma 7.2(b)). Hence, by 4.3 each color (here applied to $\gamma(v)$) appears on an external neighbor (recall, one outside of $A_{i}\cup\mathrm{All}_{i}$) of at most $4\Delta/5$ vertices in $A_{i}$.
@@ 3. $iii)$
    (Strong version of $C$) For each uncolored $A_{j}\not\in A^{\prime}$ and color $x$, at most $2\cdot\Delta^{37/40}$ vertices in $A_{j}$ have a neighbor outside of $A_{j}\cup\mathrm{All}_{j}\cup\operatorname{Big}_{j}^{+}$ that has a candidate of color $x$ or is a candidate for a node of color $x$.
@@ 7.3 Step 3: Local Matching
The remaining task – deciding on swap partners – then reduces to a local bipartite matching problem. Consider the bipartite graph with vertex bipartition $L=\mathrm{Unhappy}_{i}$ and $R=A_{i}\setminus\mathrm{Unhappy}_{i}$, with an edge between each $v\in L$ and $u\in R$ if and only if $u\in T_{v}$. By Lemma 7.6(1,2), each node in $L$ has at least $\Delta^{17/40}/40$ incident edges, while each node in $R$ has at most $\Delta^{17/40}/80$ incident edges. It follows that for every subset $S\subseteq L$, $|N(S)|\geqslant|S|$. Then, by Hall’s theorem there is a matching $M$ that saturates $L$. We perform exactly the color swaps given by the matching $M$, which yields a proper coloring by $7.$
To conclude with the proof of Lemma 4.8, it remains to argue that the $C$ property is maintained for the remaining uncolored cliques. This follows for Step 1 by Lemma 7.2(b) and for Step 3 by Lemma 7.6(3).
@@ 2. $E_{b}(i)$
    : $A_{i}$ is adjacent to a clique of $A^{\prime}$ or to an uncolored clique not in $A^{\prime}$ for which the $C$ property was broken.
@@
Let us now bound ${\mathbb{P}}[E_{b}(i)]$. Fix an arbitrary color $x$ and a clique $A_{j}$. Due to Lemma 4.1 the external degree of $A_{i}$ is bounded by $10^{8}\sqrt{\Delta}$. For a fixed external neighbor $u\in A_{\ell}\in A^{\prime\prime}$ of $A_{j}$, the probability that it receives color $x$ in the random permutation of $A_{j}$ is $1/|A_{j}|\leqslant 2/\Delta$ (by Lemma 4.1(b)). Furthermore, at most one vertex in each clique is assigned color $x$ and the random permutation for different cliques are independent. Hence, we can apply Lemma 4.4 with $Q=10^{8}\sqrt{\Delta}$ as it bounds external degrees from above (by Lemma 4.1(3,5)), and get that $A_{j}$ does not satisfy $C$ for color $x$ with probability at most $\exp(-\Delta^{1/40})$. By union bound on all colors and cliques adjacent to $A_{i}$, the event $E_{b}(i)$ occurs w.p. at most $\Delta^{3/2+1}\cdot\exp(-\Delta^{1/40})\leqslant\exp(-\Omega(\Delta^{1/40}))$.
@@
Consider the variable graph $H$ of this random process: it has a vertex (i.e., a variable) $v_{i}$ for each $A_{i}\in A^{\prime}$ and vertices $v_{i}$ and $v_{j}$ are connected iff an edge connects $A_{i}$ and $A_{j}$. The collections $\mathcal{A}$ and $\mathcal{B}$ from Lemma 3.2 are all 1-hop neighborhoods in $H$, i.e., all the $\{v_{i}\}\cup N_{H}(v_{i})$ for $A_{i}\in A^{\prime}$. Define $c_{\mathcal{B}^{\prime}}=0$ so that the set $\mathcal{B}^{\prime}$ in Lemma 3.2 includes the 1-hop neighborhood of every $v_{i}$ with $A_{i}\in A^{\prime\prime}$. This is a superset of $A^{\prime\prime}$ (the set of cliques for which we have an event in post-shattering), so it suffices to bound from above the sizes of the connected components induced by the events corresponding to sets of $\mathcal{B}^{\prime}$. Steps (S1) and (S2) with $c_{1}=2$ and $c_{2}=1$ are direct from the choice of sets $\{v_{i}\}\cup N_{H}(v_{i})$ because it has diameter at most 2 and every $v_{i}$ is included only in the sets of its neighbors. Step (S3) is direct from Algorithm 9. By Claim 7.8, each $A_{i}$ is included in $A^{\prime\prime}$ (i.e., a bad event occurred and $A_{i}$ uncolored its vertices) with probability at most $\exp(-\Omega(\Delta^{1/40}))$. The degree of a $v_{i}$ is at most $10^{9}\Delta^{3/2}$, because every vertex in $A_{i}$ may be adjacent to up to $10^{9}\sqrt{\Delta}$ vertices in other cliques. It implies that for $\Delta$ sufficiently large, we can pick a constant $c_{4}>3c_{1}(4c_{\mathcal{B}^{\prime}}+16)+c_{2}+c_{\mathcal{B}^{\prime}}+1=O(1)$ such that each set of $\mathcal{A}$ is kept in $\mathcal{A}^{\prime}$ with probability at most $\Delta(H)^{-c_{4}}$, i.e., Step (S4). Therefore Lemma 3.2 applies and, w.h.p., we can bound the size of the largest component in the dependency graph of the LLL of Algorithm 10 by some $\operatorname{\text{{\rm poly}}}(\log n)$. Let us conclude the proof by proving that the resulting coloring indeed satisfies (a) and (b). For (b), it simply follows from the fact that after retractions (Algorithm 10), the $C$ is preserved, and since events $E_{b}(i)$ are avoided in post-shattering, $C$ is also preserved by this coloring step (with a fresh budget). To see why (a) holds, observe that every $A\notin A^{\prime\prime}$ satisfies (a) after Algorithm 10 (otherwise $E_{a}(i)$ occurs and $A_{i}$ is included in $A^{\prime\prime}$). Importantly, if a monochromatic edge $\{u,v\}$ appears where $v$ was colored by Algorithm 10 and $u$ belongs to a clique of $A^{\prime}$ that did not participate in post-shattering (not in $A^{\prime\prime}$), we orient the edge from $u$ to $v$. So, the coloring of a clique $A_{j}$ that participates in the post-shattering phase can never increase the size of $\mathrm{Unhappy}_{i}$ of an adjacent clique $A_{i}$ that does not participate in the post-shattering phase. Finally, a clique $A_{i}\in A^{\prime\prime}$ colored by Algorithm 10 satisfies (a) at the end of the algorithm, as the event $E_{a}(i)$ would otherwise occur in the post-shattering LLL.
@@ 5. $B_{4}(i)$:
Fix a clique $A_{j}$ (either in $A^{\prime}$ or uncolored) and a color $x$. For a fixed external neighbor $u\in A_{\ell}\neq A_{j}$, the vertex with color $x$ in $u$’s clique (if any) gets sampled into $u$’s candidate set $S_{u}$ or samples $u$ into its candidate set with probability at most $p\leqslant\Delta^{-23/40}$. On the other hand, by assumption in Lemma 4.8, the external degrees of uncolored cliques not in $A^{\prime}$ is at most $30\Delta^{1/4}$ (recall that in this step, we do not need to guarantee that $C$ is maintained by cliques of $A^{\prime}$, in particular, cliques of $A_{H}$ do not need to verify Lemma 7.6(3)). And since the sampling is independent, it is easy to verify that (P7.1) holds with $Q=30\Delta^{1/4}$, thus that $B_{3}(i)$ occurs because of clique $A_{j}$ with probability at most $\exp(-\Delta^{1/40})$ by Lemma $4.$ By union bound on the colors and the at most $10^{8}\Delta^{3/2}$ cliques adjacent to $A_{i}$, we obtain the desired probability bound.
@@ 8 Ultrafast Coloring High-Degree Graphs
If it was not for the $C$ constraint, one could simply run the $(\deg+1)$-list-coloring algorithm of [35] to extend the coloring to $H$. Given an instance of $(\deg+1)$-list-coloring, it is not possible in general to ensure that the $C$ constraint will be maintained. Here, we use that, thanks to $\Pi$, the nodes of $H$ have lists significantly larger than their degree.
@@
Using $\Pi$ and Lemma 4.4, we can show that each of the $O(\log^{*}n)$ steps of [35] maintain the $C$ property with high probability.
@@
Both algorithms begin with a random color trial — for the sparse vertices, only some vertices try a random color. See [35, Algorithm 3]. Since every vertex of $H$ samples one color out of a list of $U\Delta^{0.22}$ while the external degree of uncolored cliques is upper bounded by $U$ (by $\Pi$), Lemma 4.4 and union bound implies that $C$ is not maintained for some clique with probability at most $n\Delta\exp(-\Delta^{1/40})<1/\operatorname{\text{{\rm poly}}}(n)$. Note that the dummy nodes do not verify $\Pi$, as their lists are of size two, but they are not adjacent to the cliques of $F$.
@@
Hence, we can apply Lemma 4.4 as before to reason that each step of MCT maintains $C$ with probability at least $1-\exp(-\Omega(\sqrt{s_{\min}}))-\Delta\exp(-\Omega(s_{\min}))\geqslant 1-1/\operatorname{\text{{\rm poly}}}(n)$.
The analysis of slack color for nodes of $H[C_{1}\cup\ldots\cup C_{q}]$ is the same as above. We emphasize that we do not need the put-aside set, Step 3 and 7 in [35, Algorithm 5], because $\Pi$ ensures that all nodes of $H$ have more than $\log^{2}n$ colors in their lists. After slack generation, [35, Algorithm 5] runs a synchronized color trial before it calls SlackColor on (most) nodes. See [35, Algorithm 7]. In this algorithm, (most of) the vertices in each $C_{i}$ receive a random color from a selected leader. As such, the probability that any given color is tried by a vertex in some $C_{i}$ is $O(1/|C_{i}|)$. Since the vertices have $\deg(v)=|L(v)|-1\geqslant U\Delta^{0.22}$, where the equality holds because of the dummy vertices and the inequality because of $\Pi$, an almost-clique $C_{i}$ contains at least $\Omega(U\Delta^{0.22})$ vertices. And thus the $C$ constraint is maintained by the SCT as every vertex tries gets some color $x$ with probability at most $O(1/|C_{i}|)\leqslant O(\frac{1}{U\Delta^{0.22}})\leqslant\frac{1}{U\Delta^{1/5}}$.